# AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistency Multispectral CT

### v1

@@ 7 Conclusion
We studied the associated CC (DD)-data model, and discovered that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix. Based on the insights, we proposed an accurate and fast image reconstruction algorithm called AFIRE respectively from the continuous, discrete and actual-use perspectives by leveraging the simplified Newton method. Under proper conditions, we further proved the convergence of the proposed algorithm AFIRE.
Furthermore, several numerical experiments were carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in geometric-inconsistency DECT for noiseless and noisy projection data in the case of mildly full scan. Through numerical comparisons, it was shown that the proposed algorithm significantly outperforms some existing methods such as NCPD, NKM, IFBP and INTRPL with respect to the accuracy and efficiency.
We also discussed the flexibility and extensibility of the proposed algorithm. It was demonstrated that the proposed algorithm is robust and flexible with regard to the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$. Moreover, for the geometric-consistency situation, it was shown that the proposed algorithm has significant advantages over the benchmark two-step DDD method.
@@ A.1 The proof of proposition 1
For any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$, $K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})$ can be rewritten as the function $H_{q}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ which is given by
@@
where $\boldsymbol{z}=[z_{1},\ldots,z_{D}]^{\boldsymbol{\mathsf{T}}}$ and $z_{d}=\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})$. It is easy to compute that
$\frac{\partial H_{q}(\boldsymbol{z})}{\partial z_{d}}=-\int_{0}^{E_{\max}}\frac{s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)}{\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)\mathrm{d}E}b_{d}(E)\mathrm{d}E.$ (33)
By 33 and the nonnegative bounded property of functions $\{b_{d}\}_{d=1}^{D}$, we have
@@
Notice that $K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})=0$ for any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$, and using 34, we obtain
$\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\left|H_{q}(\boldsymbol{z})-H_{q}(0)\right|\\ \leq\left(\sup_{\boldsymbol{z}}\|\nabla H_{q}(\boldsymbol{z})\|_{1}\right)\|\boldsymbol{z}\|_{1}\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|.$ (35)
Consequently, for any $\boldsymbol{f}\in\mathcal{F}^{D}$, by 35, we have
$\|K^{[q]}(\boldsymbol{f})\|_{\mathcal{G}}$ $=\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\Bigl{(}\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|\Bigr{)}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\|\mathcal{P}f_{d}\|_{\mathcal{G}},$
where $\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}$ denotes the Euclidean measure on the hyperplane $\boldsymbol{\omega}^{\perp}$. Since $\mathcal{P}f_{d}\in\mathcal{G}$ for each $d$, we obtain $K^{[q]}(\boldsymbol{f})\in\mathcal{G}$.
For any $\boldsymbol{h}=[h_{1},\ldots,h_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathcal{F}^{D}$, by the chain rule, we compute
$\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}h_{d}\right)\mathrm{d}E,$ (36)
@@
$\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathrm{d}E=1\quad\text{for}~{}q=1,\ldots,Q.$
Then, from 36, we have the following estimate
$\big{\|}\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}\big{\|}_{\mathcal{G}}$ $=\big{\|}\sum_{d=1}^{D}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)b_{d}(E)\mathcal{P}h_{d}\mathrm{d}E\big{\|}_{\mathcal{G}}$ (37)
$\leq\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\big{\|}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathcal{P}h_{d}\mathrm{d}E\big{\|}_{\mathcal{G}}$
$=\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\left\|\mathcal{P}h_{d}\right\|_{\mathcal{G}}$
$\leq\left(\sup_{d,E}b_{d}(E)\right)C\sum_{d=1}^{D}\|h_{d}\|_{\mathcal{F}}.$
Note that the last inequality follows readily from the result in [30, theorem 2.10], and the $C$ is a positive constant depending only on the $\alpha=-1/2$ and the domain $\Omega$. Hence, using 36 and 37, we conclude that $K^{[q]}$ is Gâteaux differentiable, and its Gâteaux derivative is computed as 10.
Furthermore, from the continuity of $\mathcal{P}$, it is easy to obtain that the mapping $\boldsymbol{f}\mapsto\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$ is continuous. By [12, proposition 3.2.15], $K^{[q]}$ is Fréchet differentiable and its Fréchet derivative is equal to $\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$. By the definition of the norm of Cartesian space, we can further prove that $\boldsymbol{K}$ is also Fréchet differentiable. ∎
@@ A.2 The proof of lemma 1
$\psi^{[q]}_{d}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})=\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})b_{d}(E)\mathrm{d}E.$
@@
$\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\|_{F}^{2}$ $=\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\big{(}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})-\Phi^{[q]}(\hat{\boldsymbol{f}};E)\big{)}b_{d}(E)\mathrm{d}E\right)^{2}$
$\leq\eta^{2}\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E\right)^{2}=\eta^{2}\|\boldsymbol{\varphi}\|_{F}^{2}.$
Subsequently, for any $\boldsymbol{z}\in\mathbb{R}^{D}$, by Cauchy–Schwarz inequality and the property of matrix norm, we have
@@
Hence, for any $\boldsymbol{f},\boldsymbol{h}\in\mathcal{F}^{D}$, based on the formula of Fréchet derivative of $\boldsymbol{K}$ and the above estimation, we obtain
$\big{\|}\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\big{\|}_{\mathcal{G}^{Q}}^{2}$ $=\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\left\|\left(\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\right)\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})\right\|_{2}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}$
@@
$=(\eta\kappa_{F}(\boldsymbol{\varphi}))^{2}\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}^{2},$
where $\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})=[\mathcal{P}h_{1}(\boldsymbol{\omega},\boldsymbol{x}),\ldots,\mathcal{P}h_{D}(\boldsymbol{\omega},\boldsymbol{x})]^{\boldsymbol{\mathsf{T}}}\in\mathbb{R}^{D}$. The desired result is obtained. ∎

### v2

@@ 6.3 Extension to the geometric-consistent situation
[…]
6.4 Validation with the basis images generated by randomly sampled Gaussian distribution
In this work, the AFIRE algorithm is developed for the geometric-inconsistent multispectral CT with a mildly full scan. In this case, the number of measurements is comparable to the total size of the employed basis images. The sufficient projection data that obtained in this case allows us to treat the basis images as entirely unknown variables, and thus can achieve an efficient reconstruction without priors on the basis images.
To further validate the statement above, we conduct additional experiments using the basis images generated by randomly sampled Gaussian distribution as truths for the water and bone basis images. Specifically, we generate 1000 random array pairs as the basis images using numpy.random with different seeds. All the other experimental settings are consistent with those described in section 5.2. After running the AFIRE algorithm for 100 iterations, we plot the scattered distributions of the metrics of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ compared to those different random truths in fig. 16. It demonstrates that the AFIRE algorithm can indeed arrive at a high-precision reconstruction even for the randomly generated images. This highlights the robustness and effectiveness of the proposed AFIRE algorithm.
@@ 7 Conclusion
We studied the associated CC (resp. DD)-data model, and discovered that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix. Based on the insights, we proposed an accurate and fast image reconstruction algorithm called AFIRE respectively from the continuous, discrete and actual-use perspectives by leveraging the simplified Newton method. Under proper conditions, we further proved the convergence of the proposed algorithm AFIRE.
Furthermore, several numerical experiments were carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in geometric-inconsistent DECT for noiseless and noisy projection data in the case of mildly full scan. Through numerical comparisons, it was shown that the proposed algorithm significantly outperforms some existing methods such as NCPD, NKM, IFBP and INTRPL with respect to the accuracy and efficiency.
We also discussed the flexibility and extensibility of the proposed algorithm. It was demonstrated that the proposed algorithm is robust and flexible with regard to the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$, and also can indeed arrive at a high-precision reconstruction even for the randomly generated images. Moreover, for the geometric-consistent situation, it was shown that the proposed algorithm has significant advantages over the benchmark two-step DDD method.
@@ A.1 The proof of proposition 1
For any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q}$, $K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})$ can be rewritten as the function $H_{q}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ which is given by
@@
where $\boldsymbol{z}=[z_{1},\ldots,z_{D}]^{\boldsymbol{\mathsf{T}}}$ and $z_{d}=\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})$. It is easy to compute that
$\frac{\partial H_{q}(\boldsymbol{z})}{\partial z_{d}}=-\int_{0}^{E_{\max}}\frac{s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)}{\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)\mathrm{d}E}b_{d}(E)~\mathrm{d}E.$ (38)
By 38 and the nonnegative bounded property of functions $\{b_{d}\}_{d=1}^{D}$, we have
@@
Notice that $K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})=0$ for any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q}$, and using 39, we obtain
$\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\left|H_{q}(\boldsymbol{z})-H_{q}(0)\right|\\ \leq\left(\sup_{\boldsymbol{z}}\|\nabla H_{q}(\boldsymbol{z})\|_{1}\right)\|\boldsymbol{z}\|_{1}\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\left|\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|.$ (40)
Consequently, for any $\boldsymbol{f}\in\mathbfcal{F}^{D}$, by 40, we have
$\|K^{[q]}(\boldsymbol{f})\|_{\mathcal{G}_{q}}$ $=\left(\int_{\mathbb{S}^{n-1}_{q}}\int_{\Omega_{q}}\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}_{q}}\int_{\Omega_{q}}\Bigl{(}\sum_{d=1}^{D}\left|\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|\Bigr{)}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}_{q}}\int_{\Omega_{q}}\sum_{d=1}^{D}\left|\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\|\mathcal{P}_{q}f_{d}\|_{\mathcal{G}_{q}},$
where $\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}$ denotes the Euclidean measure on the hyperplane $\boldsymbol{\omega}^{\perp}$. Since $\mathcal{P}_{q}f_{d}\in\mathcal{G}_{q}$ for each $d$, we obtain $K^{[q]}(\boldsymbol{f})\in\mathcal{G}_{q}$.
For any $\boldsymbol{h}=[h_{1},\ldots,h_{D}]\in\mathbfcal{F}^{D}$, by the chain rule, we compute
$\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}h_{d}\right)\mathrm{d}E,$ (41)
@@
$\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)~\mathrm{d}E=1\quad\text{for}~q=1,\ldots,Q.$
Then, from 41, we have the following estimate
$\big{\|}\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}\big{\|}_{\mathcal{G}_{q}}$ $=\big{\|}\sum_{d=1}^{D}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)b_{d}(E)\mathcal{P}_{q}h_{d}~\mathrm{d}E\big{\|}_{\mathcal{G}_{q}}$
$\leq\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\big{\|}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathcal{P}_{q}h_{d}~\mathrm{d}E\big{\|}_{\mathcal{G}_{q}}$
$=\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\left\|\mathcal{P}_{q}h_{d}\right\|_{\mathcal{G}_{q}}$
$\leq\left(\sup_{d,E}b_{d}(E)\right)C\sum_{d=1}^{D}\|h_{d}\|_{\mathcal{F}}.$ (42)
Note that the last inequality follows readily from the result in [32, theorem 2.10], and the $C$ is a positive constant depending only on $\alpha=-1/2$ and the domain $\Omega$. Hence, using 41 and section A.1, we conclude that $K^{[q]}$ is Gâteaux differentiable, and its Gâteaux derivative is computed as 10.
Furthermore, from the continuity of $\mathcal{P}_{q}$, it is easy to obtain that the mapping $\boldsymbol{f}\mapsto\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$ is continuous. By [14, proposition 3.2.15], $K^{[q]}$ is Fréchet differentiable and its Fréchet derivative is equal to $\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$. By the definition of the norm of Cartesian space, we can further prove that $\boldsymbol{K}$ is also Fréchet differentiable. ∎
@@ A.2 The proof of lemma 1
$\psi^{[q]}_{d}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})=\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})b_{d}(E)~\mathrm{d}E.$
@@
$\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-$ $\boldsymbol{\varphi}\|_{F}^{2}$
$=\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\big{(}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})-\Phi^{[q]}(\hat{\boldsymbol{f}};E)\big{)}b_{d}(E)~\mathrm{d}E\right)^{2}$
$\leq\eta^{2}\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)~\mathrm{d}E\right)^{2}=\eta^{2}\|\boldsymbol{\varphi}\|_{F}^{2}.$
Subsequently, for any $\boldsymbol{z}\in\mathbb{R}^{D}$, by Cauchy–Schwarz inequality and the property of the matrix norm, we have
@@
Hence, for any $\boldsymbol{f},\boldsymbol{h}\in\mathbfcal{F}^{D}$, based on the formula of Fréchet derivative of $\boldsymbol{K}$ and the above estimation, we obtain
$\big{\|}\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},$ $\boldsymbol{h})\big{\|}_{\mathbfcal{G}^{Q}}^{2}$
$=\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\left\|\left(\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\right)\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})\right\|_{2}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}$
@@
$=(\eta\kappa_{F}(\boldsymbol{\varphi}))^{2}\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathbfcal{G}^{Q}}^{2},$
where $\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})=[\mathcal{P}_{1}h_{1}(\boldsymbol{\omega},\boldsymbol{x}),\ldots,\mathcal{P}_{D}h_{D}(\boldsymbol{\omega},\boldsymbol{x})]^{\boldsymbol{\mathsf{T}}}\in\mathbb{R}^{D}$. The desired result is obtained. ∎